OpenStudy (anonymous):

suppose f(x)=2x-1 g(x) = 3x (f/f)(a) for a≠1/2

OpenStudy (amistre64):

really? ..... ok, remember that anything divided by itslef is equal to 1. Does that help any?

OpenStudy (anonymous):

I'm not sure what the a ≠1/2 meant

OpenStudy (amistre64):

that just means that "a" is never equal to 1/2. other wise you would get: 0/0 which is a meaningless "number"

OpenStudy (anonymous):

so the answer is just 1?

OpenStudy (amistre64):

yep.... the answer is just "1"

OpenStudy (amistre64):

f/f = 1 regardless of what f looks like.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

hmmm, not if a=1/2 =D I still dont get why they had a g(x) in there